What are the length and width of a rectangle of the length exceeds the width by 55 inches and the perimeter is 198 inches?
Answer by stanbon(75887)

What are the length and width of a rectangle of the length exceeds the width by 55 inches and the perimeter is 198 inches?
2(L+W) = 198
L = W+55
---------------------
L+W = 99
-----
W+55+W = 99
2W = 44
Width = 22 inches
------
Length = W+55 = 77 inches
--------------
Cheers,
Stan H.
-------------